Solve the equations. check the solutions.
sqrt(2y+7)+4=y

Found 2 solutions by checkley71, edjones:
Answer by checkley71(8403)

SQRT(2Y+7)+4=Y
SQRT(2Y+7)=Y-4 NOW SQUARE BOTH SIDES OF THE EQUATION.
2Y+7=Y^2-8Y+16
Y^2-8Y-2Y+16-7
Y^2-10Y+9=0
(Y-9)(Y-1)=0
Y-9=0
Y=9 ANSWER.
Y-1=0
Y=1 ANSWER.
PROOF
SQRT(2*9+7)+4=9
SQRT(18=7)=9-4
SQRT25=5
5=5

Answer by edjones(7311)

sqrt(2y+7)+4=y
sqrt(2y+7)=y-4
2y+7=y^2-8y+16 Square both sides.
y^2-10y+9=0 subtract 2y+7 from each side.
(y-9)(y-1)=0
y=9 and y=1
Check:
for x=9: sqrt(2*9+7)+4=9...sqrt(25)+4=9...5+4=9 true
for x=1: sqrt(2+7)+4=1...sqrt(9)+4=1...3+4=1 false
y=9
Ed